Suppose that a rabbit produces, on average, 30 grams of poop per day, with a standard deviation of 6. Also, suppose that we have 38 rabbits, pooping independently. Approximate the probability that the rabbits produce over 1 kilogram of poop in a single day.
Answer #1

P( > 1000/38)

= P( > 26.32)

= P(( - )/() > (26.32- )/())

= P(Z > (26.32 - 30)/(6/))

= P(Z > -3.78)

= 1 - P(Z < -3.78)

= 1 - 0 = 1